### v1

# Did Translation Models Get More Robust Without Anyone Even Noticing?

Abstract
Neural machine translation (MT) models achieve strong results across a variety of settings, but it is widely believed that they are highly sensitive to “noisy” inputs, such as spelling errors, abbreviations, and other formatting issues. In this paper, we revisit this insight in light of recent multilingual MT models and large language models (LLMs) applied to machine translation. Somewhat surprisingly, we show through controlled experiments that these models are far more robust to many kinds of noise than previous models, even when they perform similarly on clean data. This is notable because, even though LLMs have more parameters and more complex training processes than past models, none of the open ones we consider use any techniques specifically designed to encourage robustness. Next, we show that similar trends hold for social media translation experiments – LLMs are more robust to social media text. We include an analysis of the circumstances in which source correction techniques can be used to mitigate the effects of noise. Altogether, we show that robustness to many types of noise has increased.
1 Introduction
For years, the conventional wisdom has been that neural machine translation (MT) models are highly sensitive to source-side artificial and natural noise at inference time (Belinkov and Bisk, 2018). This insight has motivated many works that seek to make MT models more robust to noise through either specialized training (Ebrahimi et al., 2018; Karpukhin et al., 2019; Park et al., 2020; Vaibhav et al., 2019) or bespoke architectures (Rust et al., 2022; Salesky et al., 2021). However, MT is increasingly being performed in a different paradigm than when these analyses and architectures were created. Previously, models were mostly trained from scratch on task-specific data, often for a single language pair. Nowadays, strong results often depend on large pretrained encoder-decoder models (NLLB Team et al., 2022), instruction-tuned large language models (LLMs) like TowerLLM (Alves et al., 2024), or opaque proprietary systems like ChatGPT.111https://chat.openai.com/ These huge models may make existing robustness techniques more expensive (specialized training takes longer with larger models) or impossible (specialized architectures cannot be grafted onto an existing pretrained system). So the question is, how necessary are these robustness techniques for LLMs? Are they still needed to mitigate the brittleness of subword-level transformers, or have larger models and training sets made today’s models sufficiently robust on their own?
In this work, we investigate these questions through experiments on social media text and synthetically noised corpora. These experiments have complementary roles: social media text contains diverse noise phenomena, but isolating their effect is not straightforward because the errors are not labeled. On the other hand, synthetic errors may differ in major ways from “naturally occurring” noise, but they are interpretable and controllable, offering a way to measure noise in vitro. By evaluating on a broad spectrum of error types, we can paint a more vivid picture of what kinds of noise, and at what quantities, cause problems for MT systems.
We make the following contributions:222Our code is available at https://github.com/deep-spin/robust-mt.
We show (§3) that large pretrained models are much more robust to synthetic source-side errors than conventional single-pair NMT models (see Figure 1), even when their performance is similar on clean data. These results hold across several language pairs and varieties of noise, even though the large models lack architectural features that obviously encourage robustness to character noise.
We introduce (§3.1) a novel technique for measuring the robustness of MT models by learning a regression to predict the quality decline as a function of how noisy the source is.
We show (§4.1) that models that are robust to synthetic errors perform better at translating social media text. We investigate the relationship between synthetic robustness and performance on “real-world” noise.
We show (§5) that source correction pipelines can be an effective approach to mitigate the impact of synthetic noise without substantially worsening performance on clean data, although they are significantly less effective with stronger models, suggesting that the benefits of source correction and model robustness are not complementary. Source correction is less effective on social media data, likely because there are not enough errors to outweigh the risk of error propagation.
2 Background
2.1 Architectures for MT
The transformer.
In recent years, mainstream MT techniques have been based on the transformer (Vaswani et al., 2017), which uses multi-headed self-attention to mix information across time steps. In the original work, transformers used an encoder-decoder paradigm similar to recurrent MT models (Bahdanau et al., 2014). These models pair an encoder over the source with a decoder, an autoregressive language model that predicts target tokens one at a time. These tokens usually come from a learned subword vocabulary (Kudo, 2018; Sennrich et al., 2016). Initially, transformer MT models were trained from scratch for a single language pair on parallel data from sources such as the OPUS parallel corpus collection (Tiedemann, 2012).
Multilingual models.
Although single language pair models often perform well, they struggle in the absence of large quantities of data, making it difficult to achieve good results on low resource language pairs. This problem can be mitigated somewhat through multilingual training with systems like M2M-100 (Fan et al., 2021) and NLLB-200 (NLLB Team et al., 2022). Low resource language pairs often benefit from training data in other languages. One challenge is language imbalance – the subword vocabulary and training procedure need to be designed to allow strong performance across covered language pairs in spite of this imbalance.
LLMs for MT.
In parallel to these MT-centric developments, transformers have increasingly been used in a transfer learning set-up in which a large model is pretrained on some generic objective for which massive data is available. Such a model can then be finetuned on a particular downstream task or set of tasks. When the pretraining objective is language modeling (Radford et al., 2018), this makes it straightforward to use the model for text generation tasks such as MT. Recent generation-oriented models are often decoder-only: to use such a model for MT, the source is simply treated as part of the decoder’s context. Such models have shown some success for MT (Hendy et al., 2023). In recent times, the paradigm has shifted from traditional finetuning to instruction tuning (Sanh et al., 2022; Wei et al., 2022), in which the finetuning data is accompanied by a prompt containing an instruction. This has been shown to give models the ability to generalize to related tasks and has proven to be effective for MT (Alves et al., 2023, 2024).
2.2 Robustness to Character Noise
Character perturbations have been shown to have a large negative impact on the performance of MT models (Belinkov and Bisk, 2018). Consequently, a number of techniques have been proposed to mitigate their impact.
Robustness through training.
A common technique to increase robustness is to train MT models on examples with added source errors. Given that high-quality corpora containing authentic errors are rare, the added noise is generally synthetic (Karpukhin et al., 2019), although it can be tuned to resemble natural errors (Martucci et al., 2021; Vaibhav et al., 2019). Whether training on synthetic noise is actually helpful for becoming robust to natural errors is an open question, with various works coming to contradictory conclusions (Belinkov and Bisk, 2018; Vaibhav et al., 2019). Another hazard is that training on noise can reduce performance on clean data (Khayrallah and Koehn, 2018), so the quantity of noise needs to be chosen carefully.
Robustness through architecture.
As an alternative to specialized training techniques, robustness can be achieved with architectures other than the ubiquitous subword-level transformer. Modeling at the character or byte level (Sutskever et al., 2011; Xue et al., 2022) means that perturbations make only small changes to the sequence of tokens that the model is exposed to, whereas these same perturbations can cause a subword-level model to be exposed to completely different subword types. This may make character- and byte-level models more robust, although the evidence is mixed (Mielke et al., 2021). These models are also much slower than subword-level models because of longer sequence lengths. As an alternative, MT models can be trained on representations that are invariant to character shuffles (Belinkov and Bisk, 2018) or on visual representations of text (Salesky et al., 2021).
3 Robustness to Synthetic Noise
Model | ende | deen | enfr | fren | enko | koen | enpt | pten |
OPUS | 84.61 | 88.24 | 87.74 | 89.21 | 86.60 | 86.35 | 88.98 | 88.55 |
NLLB | 87.82 | 89.31 | 88.11 | 89.25 | 88.51 | 87.59 | 89.23 | 89.54 |
Tower | 88.21 | 89.62 | 88.79 | 89.63 | 89.33 | 88.39 | 89.83 | 89.98 |
GPT-3.5 | 88.52 | 89.64 | 88.83 | 89.45 | 89.04 | 87.98 | 89.83 | 89.81 |
LP | HF Path | Vocab | Params |
ende | opus-mt-en-de | 58.1k | 74.4M |
deen | opus-mt-de-en | 58.1k | 74.4M |
enfr | opus-mt-tc-big-en-fr | 53.0k | 232.7M |
fren | opus-mt-tc-big-fr-en | 53.0k | 232.7M |
enko | - | 32.0k* | 209.2M |
koen | - | 32.0k* | 209.2M |
enpt | opus-mt-tc-big-en-pt | 54.7k | 234.5M |
pten | - | 60.0k | 75.4M |
In our first experiments, we evaluate how various models perform in the presence of token-level synthetic errors. Although synthetic errors differ from “naturally occurring” noise, they are adjustable and function as a stress test for MT systems.
3.1 Experiments
In all of our synthetic experiments, we adopt a simple set-up: for each translation corpus, we introduce a particular type of perturbation into some percentage of the source-side tokens. We then compare performance translating this perturbed corpus to the performance on clean data. A model’s robustness can be characterized by the steepness of its decline as the noise level is increased: a flatter slope indicates that the model handles noise better.
Data.
We use four types of synthetic perturbations, each of which is a plausible error based on the mechanics of typing. For each noise type, we corrupt 10% to 100% of whitespace-delimited tokens in the FLORES-200 devtest data (NLLB Team et al., 2022). We use the following noise types:
swap: flip two adjacent characters.
chardupe: duplicate a character.
chardrop: delete a character.
key: replace a character with an adjacent one on a keyboard. We use the QWERTZ layout for German, AZERTY for French, QWERTY for English and Portuguese, and South Korean Dubeolsik for Korean.
For Korean, we used hangul-jamo333https://github.com/jonghwanhyeon/hangul-jamo to decompose hangul characters into jamo, which represent individual keystrokes, before applying perturbations.
Models.
We use models that differ in their scope (bi- or multilingual), architecture (encoder-decoder or decoder-only), and size (74M-7B parameters).
NLLB (NLLB Team et al., 2022), like OPUS, is an encoder-decoder transformer trained on parallel text. However, NLLB is a many-to-many system trained on data in 202 languages. We use the 3.3 billion parameter version.
Tower: We use the 7 billion parameter version of TowerInstruct444https://huggingface.co/Unbabel/TowerInstruct-7B-v0.1 (Alves et al., 2024), a recently-released instruction-tuned LLM derived from Llama 2 (Touvron et al., 2023). It covers translation between 10 languages.
GPT-3.5:555Specifically, we use gpt-3.5-turbo-1106. the architecture and training data of GPT-3.5 are unknown, making it difficult to draw scientific conclusions from its results. We include it because it has shown success at MT (Hendy et al., 2023) and the related GPT-4 has been shown to be able to correct some character perturbations (Cao et al., 2023).
For NLLB, Tower, and the models with a listed HF Path in Table 2, we use public checkpoints from the Hugging Face transformers library.666https://github.com/huggingface/transformers For the pten OPUS model, we use a script777https://github.com/huggingface/transformers/blob/main/src/transformers/models/marian/convert_marian_tatoeba_to_pytorch.py to convert its Marian checkpoint888https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/models/por-eng to transformers format. We directly use checkpoints from the Tatoeba Challenge (Tiedemann, 2020) and perform inference with Marian (Junczys-Dowmunt et al., 2018) for enko999https://object.pouta.csc.fi/Tatoeba-MT-models/eng-kor/opusTCv20210807-sepvoc_transformer-big_2022-07-28.zip and koen101010https://object.pouta.csc.fi/Tatoeba-MT-models/kor-eng/opusTCv20210807-sepvoc_transformer-big_2022-07-28.zip because the conversion script failed for them.
Inference.
We generate translations using beam search (Reddy et al., 1977) with a beam size of 5 for all models except GPT-3.5. For GPT-3.5, we sample with temperature 0. We use prompts for Tower111111‘‘Translate the following text from [source language] to [target language].nSource:[source text]n[target language]:’’ and GPT-3.5.121212‘‘Translate this sentence from [source language] to [target language].nSource:[source text]nTarget:’’
Evaluation.
Our base metric for scoring the translation performance on a corpus is COMET (Rei et al., 2020).131313Specifically, we use COMET-22 (Rei et al., 2022a). COMET computes a normalized score for a hypothesis , conditioned on the source and a reference . When we compute scores for translations from noisy data, we provide the COMET model the clean source, not the noisy version that was actually used to generate hypotheses. We measure the trajectory of performance as the amount of noise is increased, as depicted in Figure 1. To represent this trajectory as a single number, for each configuration we fit a linear regression to predict how much COMET declines relative to the clean performance141414There is no need to learn an intercept term because the decline is relative to the model’s clean performance. as a function of the proportion of noised tokens. We report the learned slope, which we call COMET-slope. The higher (closer to zero) the COMET-slope is, the more robust the model is. This metric can also be interpreted as the number of COMET points that would be lost if every token were corrupted.
swap
Model
ende
deen
enfr
fren
enko
koen
enpt
pten
OPUS
-73.61
-65.80
-71.02
-66.00
-74.02
-50.02
-73.35
-64.01
NLLB
-22.45
-18.20
-21.81
-20.75
-19.65
-22.89
-21.03
-20.34
Tower
-19.42
-28.54
-18.70
-27.64
-18.48
-26.16
-17.39
-28.79
GPT-3.5
-3.89
-4.36
-4.46
-5.85
-4.79
-20.89
-3.76
-6.78
chardrop
Model
ende
deen
enfr
fren
enko
koen
enpt
pten
OPUS
-69.40
-54.17
-64.33
-50.72
-71.37
-44.14
-70.13
-53.15
NLLB
-22.13
-16.51
-21.56
-17.10
-18.33
-20.81
-20.89
-18.52
Tower
-18.42
-19.34
-18.64
-17.93
-15.19
-24.62
-17.89
-20.80
GPT-3.5
-6.59
-6.55
-7.32
-5.68
-6.72
-17.81
-6.63
-7.09
chardupe
Model
ende
deen
enfr
fren
enko
koen
enpt
pten
OPUS
-56.35
-35.17
-47.42
-33.00
-65.89
-31.11
-58.98
-35.65
NLLB
-3.73
-4.02
-3.60
-5.36
-2.82
-4.19
-3.76
-4.27
Tower
-3.21
-4.13
-2.88
-4.39
-3.70
-7.66
-2.79
-3.76
GPT-3.5
-1.14
-1.36
-1.32
-1.42
-1.42
-5.64
-0.98
-1.44
key
Model
ende
deen
enfr
fren
enko
koen
enpt
pten
OPUS
-74.13
-64.46
-73.35
-66.05
-76.68
-38.50
-76.20
-66.02
NLLB
-26.20
-20.17
-25.07
-21.80
-23.69
-26.41
-24.25
-23.54
Tower
-28.64
-29.49
-27.13
-31.89
-25.89
-23.80
-26.33
-36.55
GPT-3.5
-8.19
-9.17
-8.17
-8.63
-8.91
-16.31
-7.78
-10.27
Results.
The results in Table 1 show that on clean data, Tower and GPT-3.5 are the two strongest systems, with at least one of them recording the highest COMET score for all eight language pairs. The gap between the strongest system and the much smaller OPUS models is at most COMET points for all language pairs except ende. However, Table 3 shows that the differences become much larger on noisy data. For all language pairs and noise types, OPUS suffers far more from perturbations than any of the other models do. On the other end of the spectrum, GPT-3.5 is almost always more robust than other models, sometimes strikingly so. NLLB and Tower are between these two extremes – they are far more robust than OPUS, but only come close to GPT-3.5 for chardupe noise. For swap and chardrop noise, NLLB is more robust than Tower when translating to English, while the reverse is true when translating from English. This trend is less consistent for chardupe noise. For key noise, NLLB is more robust than Tower for every pair except koen.
3.2 Analysis
Fluency and adequacy.
All models decline as the noise level is increased, but they do not decline in the same way. In Figure 2 we show how the fluency of the model’s predictions, measured by the perplexity of GPT-2 (Radford et al., 2019), is related to the quality of those predictions in terms of COMET. For fren swaps, it is clear that NLLB and GPT-3.5 continue producing fluent English text even as the noise level increases.
Size and multilinguality.
From these experiments, it is tempting to conclude that robustness depends largely on model size (OPUS is times smaller than any other system) or on multilinguality (all systems except OPUS are multilingual). However, Figure 3 tells a different story. We reran swap noise experiments with three extra models: the 600M and 1.3 billion parameter versions of NLLB (the former of which is distilled from the 54B version of NLLB) and the 1.2 billion parameter version of M2M (Fan et al., 2021). Despite NLLB-1.3B and M2M-1.2B being multilingual models of similar sizes, they do not respond the same to noise: NLLB-1.3B follows a similar curve to NLLB-3.3B, while M2M-1.2B suffers nearly as much as OPUS. NLLB-600M is somewhere between these extremes. Further work is needed to determine what factors influence this kind of robustness.
Tokenizer robustness.
Introducing perturbations affects not only translation quality but also runtime. Perturbations create character sequences that are less similar to the data that tokenizers are trained on, which leads to more pieces being used to encode the sentence. This is true even for chardrop noise, which increases the length of the tokenized sequence even as it shortens the detokenized sequence. In Table 4, we compare tokenizers by their fertility — the average number of subword pieces per whitespace word — on clean and key data. While OPUS tokenizers generally have very low fertility on clean data, it increases more than the other tokenizers, suggesting the tokenizer itself is less robust to character perturbations. It is also notable that Tower and GPT-3.5 have high fertility even on clean Korean text. While this is a symptom of tokenizer unfairness in large models (Petrov et al., 2023), it can also be a sign of tokenizer robustness: the higher the fertility, the closer the model is to byte-level tokenization. This results in noisy token sequences that are much closer to the clean sequences for Tower and GPT-3.5, as can be seen in terms of F1 in Table 5. The same trend does not hold for the other languages.
English | Portuguese | Korean | ||||
clean | key | clean | key | clean | key | |
OPUS | 1.25 | 2.71 | 1.29 | 2.62 | 1.75 | 2.93 |
NLLB | 1.40 | 2.20 | 1.53 | 2.21 | 2.03 | 3.02 |
Tower | 1.42 | 2.47 | 1.88 | 2.55 | 6.37 | 7.44 |
GPT-3.5 | 1.24 | 2.23 | 1.71 | 2.34 | 4.17 | 5.12 |
swap | chardrop | chardupe | key | |
OPUS | 21.6 | 27.3 | 36.9 | 33.8 |
NLLB | 27.6 | 35.1 | 45.7 | 42.3 |
Tower | 50.0 | 62.4 | 74.8 | 71.9 |
GPT-3.5 | 39.5 | 52.3 | 65.5 | 62.8 |
4 Robustness to Social Media Text
ende | deen | enes | esen | |||||||||
Model | FB | FC | QE | FB | FC | QE | FB | FC | QE | FB | FC | QE |
OPUS | 82.1 | 81.81 | 6.54 | 75.4 | 88.44 | 4.04 | 85.4 | 84.23 | 4.51 | 86.2 | 88.28 | 2.80 |
NLLB | 86.1 | 86.86 | 2.73 | 75.8 | 87.54 | 2.70 | 88.2 | 88.45 | 2.16 | 88.0 | 89.11 | 2.39 |
Tower | 87.3 | 88.65 | 1.74 | 74.6 | 89.85 | 2.18 | 87.9 | 89.25 | 1.50 | 85.5 | 88.96 | 1.98 |
GPT-3.5 | 87.1 | 88.45 | 1.15 | 83.1 | 91.86 | 1.32 | 89.1 | 90.14 | 0.72 | 87.9 | 91.23 | 0.91 |
The previous experiments show that large translation models and LLMs are more robust to synthetic character perturbations than conventional MT models. But is this result applicable to “authentically noisy” domains such as social media text? The nature of “noise” here is different than in the synthetic task: social media text does not necessarily contain many errors (Rello and Baeza-Yates, 2012), but the domain is very different from FLORES. These factors make it difficult to isolate the effect of noise from the general domain adaptation problem. In an ideal world, we would have a translation corpus in which each example is a triple consisting of an original noisy source sequence, a manually annotated cleaned source sequence, and a gold standard translation. This would allow translations of clean and noisy versions of the same source to be compared on some reference-based metric, isolating the effect of the errors. Unfortunately, to our knowledge no such corpus exists, so instead we perform two complementary investigations. First, we evaluate our models on MTNT (Michel and Neubig, 2018), a noisy social media translation corpus. Although this is a useful test of our models’ capabilities, the noise in the corpus is not labeled and there is no clean version of the same data to compare to. This motivates our second experiment, in which we translate data from MultiLexNorm (van der Goot et al., 2021), a lexical normalization benchmark. Together, these two experiments allow us to see both which models succeed and how badly they fail.
4.1 MTNT Experiments
MTNT pairs reddit posts with high-quality professional translations. Although the references are somewhat clean, the sources are only lightly filtered, making them potentially noisy. Unfortunately no cleaned sources exist, making the effect of noise difficult to isolate. Despite this difficulty, it is often used as a robustness benchmark (Karpukhin et al., 2019; Park et al., 2020; Salesky et al., 2021; Vaibhav et al., 2019, inter alia).
Connection to synthetic noise.
Michel and Neubig (2018) estimate that percent of English tokens in MTNT enfr are misspelled. This is much less noise than we used in our synthetic experiments, but is higher than in formal corpora.
Finetuning.
We finetuned OPUS and NLLB on the MTNT train sets for enfr, using early stopping with patience 3 and validating by loss every 100 steps. Other hyperparameters are in Table 7.
OPUS | NLLB | |
Initial learning rate | ||
Batch size | ||
Batch type | Tokens | Sequences |
#A6000 GPUs | 1 | 4 |
Max training | 20k steps | 3 epochs |
Method | enfr | fren |
OPUS | 77.51 | 80.07 |
+finetune | 79.37 | 81.67 |
NLLB | 79.20 | 80.70 |
+finetune | 80.85 | 82.60 |
Tower | 80.65 | 82.80 |
ChatGPT | 81.33 | 84.72 |
Results.
Results are shown in Table 8. GPT-3.5 performs best. Among other models, Tower is generally strongest, although NLLB roughly equals it after finetuning. OPUS performs worst.
4.2 MultiLexNorm Experiments
While MTNT is an established benchmark and useful sanity check, it is not controllable like our synthetic experiments; we cannot isolate the effect of noise because there is no non-noisy version of the corpus. Therefore we pivot to evaluate models on translating MultiLexNorm (van der Goot et al., 2021), a lexical normalization dataset that pairs social media text primarily from Twitter with manually cleaned versions of the same. Switching from MTNT to MultiLexNorm comes with a trade-off: in order to gain clean sources, we lose references.
Data.
The MultiLexNorm dataset covers 13 languages, but many of the corpora are tiny or are in languages that Tower does not cover, so we only use English, German, and Spanish151515For experiments involving Spanish, we use Helsinki-NLP/opus-mt-tc-big-en-es (234.8M parameters, 55k vocabulary) and Helsinki-NLP/opus-mt-es-en (77.9M parameters, 65k vocabulary) from transformers. as sources. In experiments with English sources, we translate to German and Spanish; otherwise, we translate to English. Statistics are presented in Table 9.
Evaluation.
As MultiLexNorm lacks reference translations, we use three reference-free evaluation techniques. First, we use faux-BLEU (Anastasopoulos, 2019), which computes (Papineni et al., 2002), where is the hypothesis computed from the noisy source, is the hypothesis computed from the clean source, and is treated as a pseudoreference.161616Specifically, we use spBLEU (NLLB Team et al., 2022). By analogy we also compute faux-COMET. These faux-metrics measure the similarity between and , with faux-BLEU being a lexical metric that captures surface-level features, while faux-COMET is more semantic. In addition, we use a new metric that we dub QE. Given , , noisy and clean source sequences and , and a reference-free quality estimation metric QE, . We use COMETKiwi (Rei et al., 2022b) to compute QE. A QE close to zero means that a model produces similar-quality outputs for both inputs, indicating robustness, whereas a large positive value indicates that translation quality suffers on noisy data.
Results.
Table 6 shows the performance of all models with both noisy and gold-standard cleaned versions of the corpora. In terms of QE, GPT-3.5 performs best for all language pairs. It also has the best faux-COMET for all except ende (where Tower passes it). By faux-BLEU it is outperformed by Tower for ende and NLLB for esen. There is a contrast between the performance of NLLB and Tower. While Tower has the better QE for all language pairs and the better faux-COMET for all except esen, NLLB outperforms it by faux-BLEU for all except ende. This suggests that NLLB preserves lexical structure, while Tower preserves “deeper” features.
5 Source Correction
Initial learning rate | |
Batch size | |
Max train steps | k |
Validation steps | |
Patience |
So far we have shown that OPUS is less robust to synthetic noise than larger models and performs worse on social media text. Next we handle noise with source correction pipelines. Pipelines cannot be learned end-to-end and introduce a risk of error propagation, but they are modular and interpretable, and can be used with closed-source models.
5.1 Correction Models
swap | chardrop | chardupe | key | |
None | 40.0 | 47.0 | 68.2 | 45.7 |
JamSpell | 75.5 | 63.5 | 91.7 | 78.0 |
mT5-Large | 85.8 | 83.2 | 95.0 | 87.9 |
We use two architectures for source correction:
JamSpell (Ozinov, 2019) is a non-neural spell-checker with trigram context. We use the default English model from the package.171717https://github.com/bakwc/JamSpell
We finetune mT5-based (Xue et al., 2021) correctors on the Prob+Word dataset from NeuSpell181818https://github.com/neuspell/neuspell (Jayanthi et al., 2020). We use the mT5-Large (1.2B parameter) version of the architecture with the hyperparameter ranges in Table 10. We select the checkpoint with the best chrF (Popović, 2015) when decoding a noised version of the FLORES English dev set. We decode with a beam size of 5.
These two models are minimalistic and maximalistic approaches to correction. JamSpell is fast (more than 10k tokens per second on a laptop CPU) but uses only trigram context. On the other hand, mT5 is a large encoder-decoder that leverages a long context. Its subword vocabulary allows it to make open-vocabulary corrections, but it is much slower than JamSpell and is prone to hallucination.
5.2 Correcting Synthetic Errors
Model | ende | enfr | enko | enpt |
OPUS | -0.44 | -0.45 | -0.30 | -0.38 |
NLLB | -0.33 | -0.39 | -0.16 | -0.37 |
Tower | -0.28 | -0.30 | 0.00 | -0.26 |
GPT-3.5 | -0.30 | 0.32 | -0.10 | -0.32 |
Model | ende | enfr | enko | enpt |
OPUS | 0.10 | 0.00 | 0.01 | -0.01 |
NLLB | -0.04 | -0.03 | -0.03 | 0.03 |
Tower | -0.04 | -0.08 | -0.03 | -0.07 |
GPT-3.5 | -0.11 | -0.08 | 0.07 | -0.04 |
MT results.
The effect of correcting synthetic enfr errors is shown in Figure 4. For OPUS models, source correction is often beneficial, with gains of more than 40 COMET points for mT5 over the baseline when at least 50% of tokens are noised. Source correction brings smaller performance gains for NLLB and Tower. Source correction has a slight negative effect on GPT-3.5, indicating that extremely robust models have little to gain from source correction. However, Table 12 shows that correcting clean inputs is nearly harmless. mT5 never loses more than COMET points. Even the weaker JamSpell never causes a degradation of more than half a COMET point, suggesting that it could be useful because of its fast runtime.
5.3 Correcting MTNT
Method | OPUS | NLLB | Tower | GPT-3.5 |
Base | 77.51 | 79.20 | 80.65 | 81.33 |
Finetuned | 79.37 | 80.85 | - | - |
JamSpell | 76.42 | 78.09 | 79.17 | 79.68 |
mT5 | 76.41 | 77.88 | 79.34 | 80.16 |
Oracle | 78.65 | 80.11 | 81.59 | 82.16 |
Results.
Social media data is a challenge for both of our source correctors because they were trained on other domains. Using either corrected source leads to a decline of at least one COMET point compared to the raw source. Although an oracle delivers some gains, it still cannot equal the performance of finetuning when it is available. These results seem to confirm that MTNT is not particularly noisy in terms of spelling errors (Karpukhin et al., 2019; Michel and Neubig, 2018).
Per-sentence analysis.
Based on the results in Table 13, it would be easy to conclude that source correction does not work for MTNT. However, Figure 5 shows that some sentences do benefit from correction, and others are not harmed. In these terms, mT5 appears to be a higher-risk corrector than JamSpell, improving translation quality for of examples (versus for JamSpell) while making quality worse for of examples (versus for JamSpell). This, in addition to the oracle performance, suggests that further research could make source correction a useful tool in this domain.
6 Conclusion
We presented several experiments testing the robustness of MT systems to synthetic and natural noise. On synthetic noise, we showed that large multilingual MT models and LLMs are far more robust than older techniques. The experiments on social media translation showed that the same larger models also worked better on natural noise. We added further support for this conclusion through reference-free translation experiments with a novel evaluation metric based on quality estimation. Finally, we exhibited circumstances in which pipeline-based source correction techniques can improve performance on noisy text, both synthetic and natural.
Acknowledgments
This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882- 00000055 (Center for Responsible AI), and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020.

### v2

@@
 Altogether, we show that robustness to many types of noise has increased.
+Did Translation Models Get More Robust Without Anyone Even Noticing?
+Ben Peters and André F.
+T.
+Martins Instituto de Telecomunicações, Lisbon, Portugal Instituto Superior Técnico, Universidade de Lisboa, Lisbon, Portugal ELLIS Unit Lisbon (LUMLIS), Lisbon, Portugal Unbabel, Lisbon, Portugal benzurdopeters@gmail.com, andre.t.martins@tecnico.ulisboa.pt
 1 Introduction
@@
 However, MT is increasingly being performed in a different paradigm than when these analyses and architectures were created.
-Previously, models were mostly trained from scratch on task-specific data, often for a single language pair.
-Nowadays, strong results often depend on large pretrained encoder-decoder models (NLLB Team et al., 2022), instruction-tuned large language models (LLMs) like TowerLLM (Alves et al., 2024), or opaque proprietary systems like ChatGPT.111https://chat.openai.com/ These huge models may make existing robustness techniques more expensive (specialized training takes longer with larger models) or impossible (specialized architectures cannot be grafted onto an existing pretrained system).
-So the question is, how necessary are these robustness techniques for LLMs?
-Are they still needed to mitigate the brittleness of subword-level transformers, or have larger models and training sets made today’s models sufficiently robust on their own?
+Previously, models were mostly trained from scratch on task-specific data.
+Nowadays, strong results often depend on instruction-tuned large language models (LLMs) like TowerLLM (Alves et al., 2024) or opaque proprietary systems like ChatGPT.111https://chat.openai.com/ These huge models may make existing robustness techniques more expensive (due to higher parameter counts) or impossible (specialized architectures cannot be grafted onto an existing pretrained system).
+So the question is, are these robustness techniques still necessary in the era of LLMs, or have larger models and training sets made today’s models sufficiently robust on their own?
 In this work, we investigate these questions through experiments on social media text and synthetically noised corpora.
-These experiments have complementary roles: social media text contains diverse noise phenomena, but isolating their effect is not straightforward because the errors are not labeled.
-On the other hand, synthetic errors may differ in major ways from “naturally occurring” noise, but they are interpretable and controllable, offering a way to measure noise in vitro.
+These experiments play complementary roles: social media text contains diverse noise phenomena, but their effect is hard to isolate because the errors are unlabeled.
+On the other hand, synthetic errors differ from real-world noise, but they are interpretable and controllable, offering a way to measure noise in vitro.
 By evaluating on a broad spectrum of error types, we can paint a more vivid picture of what kinds of noise, and at what quantities, cause problems for MT systems.
-We make the following contributions:222Our code is available at https://github.com/deep-spin/robust-mt.
-We show (§3) that large pretrained models are much more robust to synthetic source-side errors than conventional single-pair NMT models (see Figure 1), even when their performance is similar on clean data.
-These results hold across several language pairs and varieties of noise, even though the large models lack architectural features that obviously encourage robustness to character noise.
+We make the following contributions:222Our code is available at https://github.com/utter-project/robust-mt.
+We show (§3) that large pretrained models are much more robust to synthetic errors than conventional NMT models (see Figure 1), even when they perform similarly on clean data.
+This result holds across noise types and language pairs, even though the large models lack architectural features that specifically encourage robustness to character noise.
 We introduce (§3.1) a novel technique for measuring the robustness of MT models by learning a regression to predict the quality decline as a function of how noisy the source is.
@@
 We investigate the relationship between synthetic robustness and performance on “real-world” noise.
-We show (§5) that source correction pipelines can be an effective approach to mitigate the impact of synthetic noise without substantially worsening performance on clean data, although they are significantly less effective with stronger models, suggesting that the benefits of source correction and model robustness are not complementary.
-Source correction is less effective on social media data, likely because there are not enough errors to outweigh the risk of error propagation.
+We show (§5) that finetuning on noisy translation data and source correction pipelines are both effective approaches to mitigate synthetic noise without substantially worsening performance on clean data, allowing conventional NMT models to become more robust than GPT-3.5 to 3 out of 4 synthetic noise types.
+Combining correction with 7-13B parameter LLM-based translation models yields even higher robustness, allowing these pipelines to surpass GPT-3.5 on all of our synthetic benchmarks, often by a wide margin.
+Although correction is less effective for social media data on the whole, many individual examples benefit from it, suggesting that identifying these examples is a future direction.
 2 Background
@@
 These models pair an encoder over the source with a decoder, an autoregressive language model that predicts target tokens one at a time.
-These tokens usually come from a learned subword vocabulary (Kudo, 2018; Sennrich et al., 2016).
+These tokens usually come from a subword vocabulary (Kudo, 2018; Sennrich et al., 2016).
 Initially, transformer MT models were trained from scratch for a single language pair on parallel data from sources such as the OPUS parallel corpus collection (Tiedemann, 2012).
 Multilingual models.
-Although single language pair models often perform well, they struggle in the absence of large quantities of data, making it difficult to achieve good results on low resource language pairs.
-This problem can be mitigated somewhat through multilingual training with systems like M2M-100 (Fan et al., 2021) and NLLB-200 (NLLB Team et al., 2022).
+Although single language pair models often perform well, they struggle in the absence of large quantities of data, making it difficult to achieve good results in low resource language settings.
+This problem can be mitigated through multilingual training with systems like M2M-100 (Fan et al., 2021) and NLLB-200 (NLLB Team et al., 2022).
 Low resource language pairs often benefit from training data in other languages.
@@
 LLMs for MT.
-In parallel to these MT-centric developments, transformers have increasingly been used in a transfer learning set-up in which a large model is pretrained on some generic objective for which massive data is available.
-Such a model can then be finetuned on a particular downstream task or set of tasks.
-When the pretraining objective is language modeling (Radford et al., 2018), this makes it straightforward to use the model for text generation tasks such as MT.
-Recent generation-oriented models are often decoder-only: to use such a model for MT, the source is simply treated as part of the decoder’s context.
-Such models have shown some success for MT (Hendy et al., 2023).
-In recent times, the paradigm has shifted from traditional finetuning to instruction tuning (Sanh et al., 2022; Wei et al., 2022), in which the finetuning data is accompanied by a prompt containing an instruction.
-This has been shown to give models the ability to generalize to related tasks and has proven to be effective for MT (Alves et al., 2023, 2024).
+In parallel to these MT-centric developments, transformers have increasingly been used in a transfer learning set-up in which a model is pretrained on a generic objective for which massive data is available.
+The model can then be finetuned on one or more downstream tasks.
+When the pretraining objective is language modeling (Radford et al., 2018), it is straightforward to use the model for generation tasks, including MT (Hendy et al., 2023).
+Recently, the paradigm has shifted from traditional finetuning to instruction tuning (Sanh et al., 2022; Wei et al., 2022), in which the finetuning data is accompanied by an instructional prompt.
+This has been shown to give models the ability to generalize to related tasks and has proven effective for MT (Alves et al., 2023, 2024).
 2.2 Robustness to Character Noise
-Character perturbations have been shown to have a large negative impact on the performance of MT models (Belinkov and Bisk, 2018).
-Consequently, a number of techniques have been proposed to mitigate their impact.
+Character perturbations can have a large negative impact on MT model performance (Belinkov and Bisk, 2018). Consequently, a number of techniques have been proposed to mitigate their impact.
 Robustness through training.
@@
 Whether training on synthetic noise is actually helpful for becoming robust to natural errors is an open question, with various works coming to contradictory conclusions (Belinkov and Bisk, 2018; Vaibhav et al., 2019).
-Another hazard is that training on noise can reduce performance on clean data (Khayrallah and Koehn, 2018), so the quantity of noise needs to be chosen carefully.
 Robustness through architecture.
@@
 3 Robustness to Synthetic Noise
-Model | ende | deen | enfr | fren | enko | koen | enpt | pten |
-OPUS | 84.61 | 88.24 | 87.74 | 89.21 | 86.60 | 86.35 | 88.98 | 88.55 |
-NLLB | 87.82 | 89.31 | 88.11 | 89.25 | 88.51 | 87.59 | 89.23 | 89.54 |
-Tower | 88.21 | 89.62 | 88.79 | 89.63 | 89.33 | 88.39 | 89.83 | 89.98 |
-GPT-3.5 | 88.52 | 89.64 | 88.83 | 89.45 | 89.04 | 87.98 | 89.83 | 89.81 |
-LP | HF Path | Vocab | Params |
-ende | opus-mt-en-de | 58.1k | 74.4M |
-deen | opus-mt-de-en | 58.1k | 74.4M |
-enfr | opus-mt-tc-big-en-fr | 53.0k | 232.7M |
-fren | opus-mt-tc-big-fr-en | 53.0k | 232.7M |
-enko | - | 32.0k* | 209.2M |
-koen | - | 32.0k* | 209.2M |
-enpt | opus-mt-tc-big-en-pt | 54.7k | 234.5M |
-pten | - | 60.0k | 75.4M |
-In our first experiments, we evaluate how various models perform in the presence of token-level synthetic errors.
-Although synthetic errors differ from “naturally occurring” noise, they are adjustable and function as a stress test for MT systems.
+xxen
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+88.17
+89.19
+86.35
+88.36
+88.02
+NLLB
+89.28
+89.29
+87.69
+89.72
+89.00
+TI
+89.77
+89.76
+88.69
+90.16
+89.60
+GPT-3.5
+89.64
+89.45
+87.98
+89.81
+89.22
+enxx
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+84.02
+87.63
+86.58
+88.94
+86.79
+NLLB
+88.07
+88.30
+88.48
+89.58
+88.61
+TI
+88.57
+89.16
+90.12
+90.02
+89.47
+GPT-3.5
+88.52
+88.83
+89.04
+89.83
+89.05
+In our first experiments, we evaluate how models perform in the presence of token-level synthetic errors.
+Although these errors differ from “naturally occurring” noise, they are adjustable and function as a stress test for MT systems.
 3.1 Experiments
@@
 We use four types of synthetic perturbations, each of which is a plausible error based on the mechanics of typing.
-For each noise type, we corrupt 10% to 100% of whitespace-delimited tokens in the FLORES-200 devtest data (NLLB Team et al., 2022).
+For each noise type, we create ten noised versions of the FLORES-200 devtest set (NLLB Team et al., 2022) corresponding to noise levels .
+Within a version of the corpus, each whitespace-delimited token is perturbed with probability and otherwise not altered.
+Therefore each token can be perturbed at most once.
 We use the following noise types:
 swap: flip two adjacent characters.
-chardupe: duplicate a character.
-chardrop: delete a character.
-key: replace a character with an adjacent one on a keyboard. We use the QWERTZ layout for German, AZERTY for French, QWERTY for English and Portuguese, and South Korean Dubeolsik for Korean.
-For Korean, we used hangul-jamo333https://github.com/jonghwanhyeon/hangul-jamo to decompose hangul characters into jamo, which represent individual keystrokes, before applying perturbations.
+dupe: duplicate a character.
+drop: delete a character.
+key: replace a character with an adjacent key. Further details are in Appendix A.
 Models.
-We use models that differ in their scope (bi- or multilingual), architecture (encoder-decoder or decoder-only), and size (74M-7B parameters).
+We use models that differ in their scope (bi- or multilingual), architecture (encoder-decoder or decoder-only), and size (74M-13B parameters).
 NLLB (NLLB Team et al., 2022), like OPUS, is an encoder-decoder transformer trained on parallel text.
@@
 We use the 3.3 billion parameter version.
-Tower: We use the 7 billion parameter version of TowerInstruct444https://huggingface.co/Unbabel/TowerInstruct-7B-v0.1 (Alves et al., 2024), a recently-released instruction-tuned LLM derived from Llama 2 (Touvron et al., 2023).
-It covers translation between 10 languages.
-GPT-3.5:555Specifically, we use gpt-3.5-turbo-1106.
-the architecture and training data of GPT-3.5 are unknown, making it difficult to draw scientific conclusions from its results.
-We include it because it has shown success at MT (Hendy et al., 2023) and the related GPT-4 has been shown to be able to correct some character perturbations (Cao et al., 2023).
-For NLLB, Tower, and the models with a listed HF Path in Table 2, we use public checkpoints from the Hugging Face transformers library.666https://github.com/huggingface/transformers For the pten OPUS model, we use a script777https://github.com/huggingface/transformers/blob/main/src/transformers/models/marian/convert_marian_tatoeba_to_pytorch.py to convert its Marian checkpoint888https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/models/por-eng to transformers format.
-We directly use checkpoints from the Tatoeba Challenge (Tiedemann, 2020) and perform inference with Marian (Junczys-Dowmunt et al., 2018) for enko999https://object.pouta.csc.fi/Tatoeba-MT-models/eng-kor/opusTCv20210807-sepvoc_transformer-big_2022-07-28.zip and koen101010https://object.pouta.csc.fi/Tatoeba-MT-models/kor-eng/opusTCv20210807-sepvoc_transformer-big_2022-07-28.zip because the conversion script failed for them.
+TI: We use the 13 billion parameter version of TowerInstruct-v0.1 (Alves et al., 2024), an instruction-tuned LLM that can translate between 10 languages.
+As TI and GPT-3.5 were both trained on closed data, it is possible that they were trained on our test sets.
+We include them because of the lack of high-quality fully-open LLM-based translation systems.
+swap
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-65.05
+-65.41
+-36.18
+-63.44
+-57.52
+NLLB
+-18.14
+-20.97
+-23.79
+-20.81
+-20.93
+TI
+-27.61
+-27.01
+-23.45
+-25.54
+-25.90
+GPT-3.5
+-4.36
+-5.85
+-20.89
+-6.78
+-9.47
+drop
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-53.61
+-49.92
+-29.54
+-52.48
+-46.39
+NLLB
+-16.50
+-17.27
+-21.11
+-19.04
+-18.48
+TI
+-19.62
+-18.01
+-17.33
+-17.71
+-18.17
+GPT-3.5
+-6.55
+-5.68
+-17.81
+-7.09
+-9.28
+dupe
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-34.31
+-31.83
+-6.92
+-34.58
+-26.91
+NLLB
+-4.07
+-5.31
+-4.36
+-4.58
+-4.58
+TI
+-3.37
+-4.54
+-1.82
+-3.62
+-3.38
+GPT-3.5
+-1.36
+-1.42
+-5.64
+-1.44
+-2.47
+key
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-63.78
+-65.40
+-38.48
+-65.50
+-58.29
+NLLB
+-20.66
+-21.86
+-28.01
+-23.60
+-23.53
+TI
+-29.15
+-32.18
+-19.80
+-34.15
+-28.82
+GPT-3.5
+-9.17
+-8.63
+-16.31
+-10.27
+-11.09
+swap
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-72.01
+-69.59
+-73.99
+-72.97
+-72.14
+NLLB
+-23.33
+-22.75
+-19.68
+-22.41
+-22.04
+TI
+-16.64
+-13.71
+-12.63
+-13.44
+-14.11
+GPT-3.5
+-3.89
+-4.46
+-4.79
+-3.76
+-4.23
+drop
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-67.77
+-63.30
+-71.31
+-69.66
+-68.01
+NLLB
+-22.65
+-22.23
+-18.45
+-21.71
+-21.26
+TI
+-17.22
+-15.00
+-9.08
+-14.68
+-14.00
+GPT-3.5
+-6.59
+-7.32
+-6.72
+-6.63
+-6.81
+dupe
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-54.90
+-46.25
+-65.86
+-57.94
+-56.24
+NLLB
+-4.04
+-3.81
+-2.79
+-4.13
+-3.69
+TI
+-2.40
+-1.79
+-1.39
+-1.89
+-1.87
+GPT-3.5
+-1.14
+-1.32
+-1.42
+-0.98
+-1.21
+key
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-72.46
+-72.01
+-76.64
+-75.81
+-74.23
+NLLB
+-27.32
+-25.91
+-23.90
+-25.57
+-25.67
+TI
+-24.51
+-21.10
+-15.95
+-22.04
+-20.90
+GPT-3.5
+-8.19
+-8.17
+-8.91
+-7.78
+-8.26
 Inference.
-We generate translations using beam search (Reddy et al., 1977) with a beam size of 5 for all models except GPT-3.5.
-For GPT-3.5, we sample with temperature 0.
-We use prompts for Tower111111‘‘Translate the following text from [source language] to [target language].nSource:[source text]n[target language]:’’ and GPT-3.5.121212‘‘Translate this sentence from [source language] to [target language].nSource:[source text]nTarget:’’
+For GPT-3.5, we sample with temperature 0. For other models, we decode with a beam size of 5. Details are in Appendix C.
 Evaluation.
-Our base metric for scoring the translation performance on a corpus is COMET (Rei et al., 2020).131313Specifically, we use COMET-22 (Rei et al., 2022a).
+Our base corpus-level translation metric is COMET (Rei et al., 2020).444We use Unbabel/wmt22-comet-da (Rei et al., 2022a).
 COMET computes a normalized score for a hypothesis , conditioned on the source and a reference .
@@
 We measure the trajectory of performance as the amount of noise is increased, as depicted in Figure 1.
-To represent this trajectory as a single number, for each configuration we fit a linear regression to predict how much COMET declines relative to the clean performance141414There is no need to learn an intercept term because the decline is relative to the model’s clean performance.
+To represent this trajectory as a single number, we fit a linear regression to predict the COMET decline relative to the clean score555There is no need to learn an intercept term because the decline is relative to the model’s clean performance.
 as a function of the proportion of noised tokens.
 We report the learned slope, which we call COMET-slope.
-The higher (closer to zero) the COMET-slope is, the more robust the model is.
-This metric can also be interpreted as the number of COMET points that would be lost if every token were corrupted.
-swap
-Model
-ende
-deen
-enfr
-fren
-enko
-koen
-enpt
-pten
-OPUS
--73.61
--65.80
--71.02
--66.00
--74.02
--50.02
--73.35
--64.01
-NLLB
--22.45
--18.20
--21.81
--20.75
--19.65
--22.89
--21.03
--20.34
-Tower
--19.42
--28.54
--18.70
--27.64
--18.48
--26.16
--17.39
--28.79
-GPT-3.5
--3.89
--4.36
--4.46
--5.85
--4.79
--20.89
--3.76
--6.78
-chardrop
-Model
-ende
-deen
-enfr
-fren
-enko
-koen
-enpt
-pten
-OPUS
--69.40
--54.17
--64.33
--50.72
--71.37
--44.14
--70.13
--53.15
-NLLB
--22.13
--16.51
--21.56
--17.10
--18.33
--20.81
--20.89
--18.52
-Tower
--18.42
--19.34
--18.64
--17.93
--15.19
--24.62
--17.89
--20.80
-GPT-3.5
--6.59
--6.55
--7.32
--5.68
--6.72
--17.81
--6.63
--7.09
-chardupe
-Model
-ende
-deen
-enfr
-fren
-enko
-koen
-enpt
-pten
-OPUS
--56.35
--35.17
--47.42
--33.00
--65.89
--31.11
--58.98
--35.65
-NLLB
--3.73
--4.02
--3.60
--5.36
--2.82
--4.19
--3.76
--4.27
-Tower
--3.21
--4.13
--2.88
--4.39
--3.70
--7.66
--2.79
--3.76
-GPT-3.5
--1.14
--1.36
--1.32
--1.42
--1.42
--5.64
--0.98
--1.44
-key
-Model
-ende
-deen
-enfr
-fren
-enko
-koen
-enpt
-pten
-OPUS
--74.13
--64.46
--73.35
--66.05
--76.68
--38.50
--76.20
--66.02
-NLLB
--26.20
--20.17
--25.07
--21.80
--23.69
--26.41
--24.25
--23.54
-Tower
--28.64
--29.49
--27.13
--31.89
--25.89
--23.80
--26.33
--36.55
-GPT-3.5
--8.19
--9.17
--8.17
--8.63
--8.91
--16.31
--7.78
--10.27
+A higher (closer to zero) COMET-slope indicates a more robust model.
+This metric can be interpreted as the number of COMET points that would be lost if every token were corrupted.
 Results.
-The results in Table 1 show that on clean data, Tower and GPT-3.5 are the two strongest systems, with at least one of them recording the highest COMET score for all eight language pairs.
-The gap between the strongest system and the much smaller OPUS models is at most COMET points for all language pairs except ende.
-However, Table 3 shows that the differences become much larger on noisy data.
-For all language pairs and noise types, OPUS suffers far more from perturbations than any of the other models do.
-On the other end of the spectrum, GPT-3.5 is almost always more robust than other models, sometimes strikingly so.
-NLLB and Tower are between these two extremes – they are far more robust than OPUS, but only come close to GPT-3.5 for chardupe noise.
-For swap and chardrop noise, NLLB is more robust than Tower when translating to English, while the reverse is true when translating from English.
-This trend is less consistent for chardupe noise.
-For key noise, NLLB is more robust than Tower for every pair except koen.
+Table 1 shows that on clean data, TI records the highest COMET for all eight language pairs.
+The gap between the strongest system and the much smaller OPUS models is less than COMET for all pairs except ende and enko.
+However, Tables 2 and 3 show that OPUS suffers more from perturbations than the other models do.
+On the other end of the spectrum, GPT-3.5 is almost always the most robust system.
+NLLB and TI are between these two extremes.
+For swap and drop noise, NLLB is more robust than TI when translating to English, while the reverse is true when translating from English.
+This trend is less consistent for dupe noise.
+For key noise, NLLB is more robust than TI for every pair except koen.
+BLEU and chrF results are in Appendix H.
 3.2 Analysis
-Fluency and adequacy.
-All models decline as the noise level is increased, but they do not decline in the same way.
-In Figure 2 we show how the fluency of the model’s predictions, measured by the perplexity of GPT-2 (Radford et al., 2019), is related to the quality of those predictions in terms of COMET.
-For fren swaps, it is clear that NLLB and GPT-3.5 continue producing fluent English text even as the noise level increases.
+| Model | clean | swap | drop | dupe | key |
+| OPUS | 88.94 | -72.97 | -69.66 | -57.94 | -75.81 |
+| OPUSLLM | 85.78 | -73.05 | -69.68 | -55.63 | -74.48 |
 Size and multilinguality.
-From these experiments, it is tempting to conclude that robustness depends largely on model size (OPUS is times smaller than any other system) or on multilinguality (all systems except OPUS are multilingual).
-However, Figure 3 tells a different story.
-We reran swap noise experiments with three extra models: the 600M and 1.3 billion parameter versions of NLLB (the former of which is distilled from the 54B version of NLLB) and the 1.2 billion parameter version of M2M (Fan et al., 2021).
-Despite NLLB-1.3B and M2M-1.2B being multilingual models of similar sizes, they do not respond the same to noise: NLLB-1.3B follows a similar curve to NLLB-3.3B, while M2M-1.2B suffers nearly as much as OPUS.
-NLLB-600M is somewhere between these extremes.
-Further work is needed to determine what factors influence this kind of robustness.
+From these experiments, one might conclude that robustness depends largely on model size (OPUS is times smaller than any other system) or multilinguality (all except OPUS are multilingual).
+However, Figure 2 shows that these are not the only factors.
+We reran swap experiments with NLLB-600M, NLLB-1.3B, and M2M-1.2B (Fan et al., 2021); despite the similar sizes of NLLB-1.3B and M2M-1.2B, and the fact that they are both massively multilingual, they handle noise differently: NLLB-1.3B is similar to NLLB-3.3B, while M2M-1.2B suffers as much as OPUS.
+Impact of architecture.
+Having shown that model size does not have a strong effect on robustness (at least not for NLLB), we next investigate the impact of architecture on performance.
+Is the gap between OPUS and other models primarily due to differences in training data, or is there some aspect of the LLMs’ decoder-only structure that encourages robustness?
+To investigate, we trained a 1.3B parameter666We would have preferred to train a 13B model (similar to TI), but this was impossible due to resource constraints.
+decoder-only model on the same Tatoeba Challenge data as was used by the enpt OPUS model.
+Training details are given in Appendix D.
+The performance and robustness of this model, which we dub OPUSLLM, are shown in Table 4.
+Although its performance on clean data lags COMET points behind OPUS, the COMET-slope is similar for all four noise types, suggesting that the robustness of recent models is due to their training data, not their size or architecture.
 Tokenizer robustness.
@@
 Perturbations create character sequences that are less similar to the data that tokenizers are trained on, which leads to more pieces being used to encode the sentence.
-This is true even for chardrop noise, which increases the length of the tokenized sequence even as it shortens the detokenized sequence.
+This is true even for drop noise, which increases the length of the tokenized sequence even as it shortens the detokenized sequence.
 In Table 5, we compare tokenizers by their fertility — the average number of subword pieces per whitespace word — on clean and key data.
 While OPUS tokenizers generally have very low fertility on clean data, it increases more than the other tokenizers, suggesting the tokenizer itself is less robust to character perturbations.
-It is also notable that Tower and GPT-3.5 have high fertility even on clean Korean text.
+It is also notable that TI and GPT-3.5 have high fertility even on clean Korean text.
 While this is a symptom of tokenizer unfairness in large models (Petrov et al., 2023), it can also be a sign of tokenizer robustness: the higher the fertility, the closer the model is to byte-level tokenization.
-This results in noisy token sequences that are much closer to the clean sequences for Tower and GPT-3.5, as can be seen in terms of F1 in Table 5.
+This results in noisy token sequences that are much closer to the clean sequences for TI and GPT-3.5, as can be seen in terms of F1 in Table 6.
 The same trend does not hold for the other languages.
@@
 | NLLB | 1.40 | 2.20 | 1.53 | 2.21 | 2.03 | 3.02 |
-Tower | 1.42 | 2.47 | 1.88 | 2.55 | 6.37 | 7.44 |
+| TI | 1.42 | 2.47 | 1.88 | 2.55 | 6.37 | 7.44 |
 | GPT-3.5 | 1.24 | 2.23 | 1.71 | 2.34 | 4.17 | 5.12 |
-swap | chardrop | chardupe | key | |
+| swap | drop | dupe | key | |
 | OPUS | 21.6 | 27.3 | 36.9 | 33.8 |
 | NLLB | 27.6 | 35.1 | 45.7 | 42.3 |
-Tower | 50.0 | 62.4 | 74.8 | 71.9 |
+| TI | 50.0 | 62.4 | 74.8 | 71.9 |
 | GPT-3.5 | 39.5 | 52.3 | 65.5 | 62.8 |
 4 Robustness to Social Media Text
-ende | deen | enes | esen | |||||||||
-Model | FB | FC | QE | FB | FC | QE | FB | FC | QE | FB | FC | QE |
-OPUS | 82.1 | 81.81 | 6.54 | 75.4 | 88.44 | 4.04 | 85.4 | 84.23 | 4.51 | 86.2 | 88.28 | 2.80 |
-NLLB | 86.1 | 86.86 | 2.73 | 75.8 | 87.54 | 2.70 | 88.2 | 88.45 | 2.16 | 88.0 | 89.11 | 2.39 |
-Tower | 87.3 | 88.65 | 1.74 | 74.6 | 89.85 | 2.18 | 87.9 | 89.25 | 1.50 | 85.5 | 88.96 | 1.98 |
-GPT-3.5 | 87.1 | 88.45 | 1.15 | 83.1 | 91.86 | 1.32 | 89.1 | 90.14 | 0.72 | 87.9 | 91.23 | 0.91 |
 The previous experiments show that large translation models and LLMs are more robust to synthetic character perturbations than conventional MT models.
@@
 The nature of “noise” here is different than in the synthetic task: social media text does not necessarily contain many errors (Rello and Baeza-Yates, 2012), but the domain is very different from FLORES.
-These factors make it difficult to isolate the effect of noise from the general domain adaptation problem.
-In an ideal world, we would have a translation corpus in which each example is a triple consisting of an original noisy source sequence, a manually annotated cleaned source sequence, and a gold standard translation.
+This makes it difficult to isolate the effect of noise from the general domain adaptation problem.
+Ideally, we would have a translation corpus in which each example is a triple consisting of an original noisy source, a manually annotated cleaned source, and a gold standard translation.
 This would allow translations of clean and noisy versions of the same source to be compared on some reference-based metric, isolating the effect of the errors.
-Unfortunately, to our knowledge no such corpus exists, so instead we perform two complementary investigations.
-First, we evaluate our models on MTNT (Michel and Neubig, 2018), a noisy social media translation corpus.
-Although this is a useful test of our models’ capabilities, the noise in the corpus is not labeled and there is no clean version of the same data to compare to.
+As we are aware of only one such corpus (Bawden and Sagot, 2023), we instead perform two complementary investigations.
+First, we evaluate on MTNT (Michel and Neubig, 2018), a noisy social media MT corpus.
+Although this is a useful test of our models, the noise is not labeled and there is no clean version of the same data to compare to.
 This motivates our second experiment, in which we translate data from MultiLexNorm (van der Goot et al., 2021), a lexical normalization benchmark.
-Together, these two experiments allow us to see both which models succeed and how badly they fail.
+Together, these experiments allow us to see both which models succeed and how badly they fail.
 4.1 MTNT Experiments
 MTNT pairs Reddit posts with high-quality professional translations.
-Although the references are somewhat clean, the sources are only lightly filtered, making them potentially noisy.
+Although the references are somewhat clean, the sources are only lightly filtered, making them potentially noisy.777 percent of MTNT enfr source tokens are misspelled (Michel and Neubig, 2018).
+This is a higher rate than in formal corpora, but lower than in our synthetic experiments.
 Unfortunately no cleaned sources exist, making the effect of noise difficult to isolate.
 Despite this difficulty, it is often used as a robustness benchmark (Karpukhin et al., 2019; Park et al., 2020; Salesky et al., 2021; Vaibhav et al., 2019, inter alia).
-Connection to synthetic noise.
-Michel and Neubig (2018) estimate that percent of English tokens in MTNT enfr are misspelled. This is much less noise than we used in our synthetic experiments, but is higher than in formal corpora.
 Finetuning.
-We finetuned OPUS and NLLB on the MTNT train sets for enfr, using early stopping with patience 3 and validating by loss every 100 steps. Other hyperparameters are in Table 7.
-OPUS | NLLB | |
-Initial learning rate | ||
-Batch size | ||
-Batch type | Tokens | Sequences |
-#A6000 GPUs | 1 | 4 |
-Max training | 20k steps | 3 epochs |
+We finetuned OPUS on MTNT enfr as described in Appendix E. We dub this model r/OPUS.
 | Method | enfr | fren |
-OPUS | 77.51 | 80.07 |
-+finetune | 79.37 | 81.67 |
-NLLB | 79.20 | 80.70 |
-+finetune | 80.85 | 82.60 |
-Tower | 80.65 | 82.80 |
-ChatGPT | 81.33 | 84.72 |
+| OPUS | 77.21 | 79.64 |
+| r/OPUS | 79.22 | 81.94 |
+| NLLB | 79.33 | 80.59 |
+| TI | 81.91 | 83.66 |
+| GPT-3.5 | 81.33 | 84.72 |
 Results.
-Results are shown in Table 8. GPT-3.5 performs best. Among other models, Tower is generally strongest, although NLLB roughly equals it after finetuning. OPUS performs worst.
+Results are shown in Table 7.
+Despite in-domain finetuning benefiting OPUS by more than 2 COMET points for both enfr and fren, this does not close the gap to TI and GPT-3.5.
+This suggests that TI and GPT-3.5 benefit from their massive training corpora, which likely contain large quantities of social media text.
+In contrast, the only social media text the finetuned OPUS models have seen are MTNT’s tiny training sets (36k parallel examples for enfr, 19k for fren), plus whatever is in the Tatoeba Challenge corpora.
+| ende | deen | enes | esen | |||||||||
+| Model | FB | FC | QE | FB | FC | QE | FB | FC | QE | FB | FC | QE |
+| OPUS | 81.9 | 81.50 | 6.21 | 74.8 | 88.36 | 3.95 | 80.3 | 82.54 | 4.11 | 86.2 | 88.52 | 2.54 |
+| NLLB | 87.3 | 87.28 | 2.81 | 74.8 | 87.75 | 2.88 | 88.3 | 88.62 | 2.24 | 84.2 | 88.70 | 2.30 |
+| TI | 88.8 | 89.07 | 1.86 | 75.2 | 90.57 | 2.41 | 89.0 | 90.26 | 1.19 | 85.3 | 89.32 | 2.63 |
+| GPT-3.5 | 87.1 | 88.45 | 1.15 | 83.1 | 91.86 | 1.32 | 89.1 | 90.14 | 0.72 | 87.9 | 91.23 | 0.91 |
 4.2 MultiLexNorm Experiments
@@
 Data.
-The MultiLexNorm dataset covers 13 languages, but many of the corpora are tiny or are in languages that Tower does not cover, so we only use English, German, and Spanish151515For experiments involving Spanish, we use Helsinki-NLP/opus-mt-tc-big-en-es (234.8M parameters, 55k vocabulary) and Helsinki-NLP/opus-mt-es-en (77.9M parameters, 65k vocabulary) from transformers.
-as sources.
+We use the English, German, and Spanish data from MultiLexNorm as our translation sources.
 In experiments with English sources, we translate to German and Spanish; otherwise, we translate to English.
-Statistics are presented in Table 9.
+Statistics are presented in Appendix F.
 Evaluation.
-As MultiLexNorm lacks reference translations, we use three reference-free evaluation techniques.
-First, we use faux-BLEU (Anastasopoulos, 2019), which computes (Papineni et al., 2002), where is the hypothesis computed from the noisy source, is the hypothesis computed from the clean source, and is treated as a pseudoreference.161616Specifically, we use spBLEU (NLLB Team et al., 2022).
+As MultiLexNorm lacks reference translations, we use three reference-free metrics.
+Faux-BLEU (Anastasopoulos, 2019) computes (Papineni et al., 2002; NLLB Team et al., 2022), where and are the hypotheses computed from the noisy source and the clean source, respectively.
+is treated as a pseudoreference.
 By analogy we also compute faux-COMET.
-These faux-metrics measure the similarity between and , with faux-BLEU being a lexical metric that captures surface-level features, while faux-COMET is more semantic.
-In addition, we use a new metric that we dub QE.
+These faux-metrics measure the similarity between and , with faux-BLEU capturing lexical features while faux-COMET is “deeper.” In addition, we use a new metric that we dub QE.
 Given , , noisy and clean source sequences and , and a reference-free quality estimation metric QE, .
-We use COMETKiwi (Rei et al., 2022b) to compute QE.
+We compute QE with COMETKiwi888Specifically, we use Unbabel/wmt22-cometkiwi-da.
+(Rei et al., 2022b).
 A QE close to zero means that a model produces similar-quality outputs for both inputs, indicating robustness, whereas a large positive value indicates that translation quality suffers on noisy data.
@@
 In terms of QE, GPT-3.5 performs best for all language pairs.
-It also has the best faux-COMET for all except ende (where Tower passes it).
-By faux-BLEU it is outperformed by Tower for ende and NLLB for esen.
-There is a contrast between the performance of NLLB and Tower.
-While Tower has the better QE for all language pairs and the better faux-COMET for all except esen, NLLB outperforms it by faux-BLEU for all except ende.
-This suggests that NLLB preserves lexical structure, while Tower preserves “deeper” features.
-5 Source Correction
-Initial learning rate | |
-Batch size | |
-Max train steps | k |
-Validation steps | |
-Patience |
+It also records the best faux-BLEU for all pairs except ende, on which TI and NLLB both outperform it.
+The faux-COMET results show a split, where GPT-3.5 has the highest scores for both to-English pairs but TI passes it for ende and enes.
+5 Strategies for Mitigating Noise
 So far we have shown that OPUS is less robust to synthetic noise than larger models and performs worse on social media text.
-Next we handle noise with source correction pipelines.
-Pipelines cannot be learned end-to-end and introduce a risk of error propagation, but they are modular and interpretable, and can be used with closed-source models.
-5.1 Correction Models
-swap | chardrop | chardupe | key | |
+Next we evaluate two techniques for mitigating noise: finetuning MT models on synthetically noised data and incorporating a source correction pipeline.
+These approaches have contrasting trade-offs.
+Finetuning the MT model allows the robust translation task to be learned end-to-end.
+However, this is expensive if a model is very large (TI, NLLB) and impossible if its weights are closed (GPT-3.5).
+On the other hand, pipelines are modular, allowing the same correction system to be reused with any model.
+The drawback of this modularity is that pipelines may introduce errors as well as fixing them.
+5.1 Synthetic Experiments
+As a comparison of finetuning and source correction, we focus on enpt with the same synthetically noised corpora as in Section 3.
+We kept the training data as similar as possible between finetuning and correction experiments: in both cases, we subsampled 5 million examples from the Tatoeba Challenge enpt training set.
+We noised 15% of source tokens with each of the four noise types used in Section 3, so a total of 60% of tokens were corrupted.
+For validation, we concatenated the original FLORES enpt development set to a version of the same set in which 20% of tokens have been corrupted with each noise type.
+By including clean validation data, this favors models that do not forget how to translate clean data.
+For finetuning OPUS, these noised source examples are paired with clean Portuguese targets.
+For finetuning correction models, they are paired with the original clean English sources.
+We followed the training procedure in Appendix E for both MT and correction models.
+| Model | clean | swap | drop | dupe | key |
+| OPUS | 88.94 | -72.97 | -69.66 | -57.94 | -75.81 |
+| +finetuning | 88.52 | -2.14 | -7.59 | -0.87 | -5.01 |
+| +correction | 88.36 | -2.02 | -11.81 | -0.08 | -7.20 |
+| NLLB | 89.58 | -22.41 | -21.71 | -4.13 | -25.57 |
+| +correction | 89.11 | -1.57 | -9.09 | -0.07 | -5.47 |
+| TI | 90.02 | -13.44 | -14.68 | -1.89 | -22.04 |
+| +correction | 89.57 | -0.86 | -5.59 | -0.16 | -2.99 |
+| ChatGPT | 89.83 | -3.76 | -6.63 | -0.98 | -7.78 |
+MT Finetuning
+Model
+clean
+swap
+drop
+dupe
+key
+100k
+88.28
+-12.08
+-18.96
+-4.63
+-18.65
+1m
+88.13
+-5.21
+-11.72
+-2.04
+-9.49
+5m
+88.52
+-2.14
+-7.59
+-0.87
+-5.01
+Source Correction
+Model
+clean
+swap
+drop
+dupe
+key
+100k
+88.59
+-4.12
+-15.31
+-0.11
+-10.84
+1m
+88.30
+-2.25
+-12.83
+-0.15
+-7.83
+5m
+88.36
+-2.02
+-11.81
+-0.08
+-7.20
+Source correction (SC).
+As our base model for SC, we adopted ByT5-Small (Xue et al., 2022).
+At inference time we report the results of a pipeline that pairs the corrector with a translation model.
+In intrinsic terms, this corrector manages a chrF (Popović, 2015) of at least at correcting each noise type, and over for swaps and dupes.
+Additional results are shown in Appendix G.
+Results.
+Our main results are shown in Table 9.
+For OPUS, both MT finetuning and correction greatly improve robustness to synthetic errors, with both techniques reporting flatter COMET-slopes than GPT-3.5 for all noise types except drop.
+At a more granular level, in Figure 3 we show OPUS’s COMET scores at all levels of swap noise for MT finetuning and SC.
+It is clear that MT finetuning outperforms SC across almost all noise levels.
+Despite NLLB and TI being more robust than OPUS in general, they too benefit from SC, suggesting that its effect is complementary with the models’ inherent robustness.
+Both MT finetuning and SC slightly degrade model performance on clean translation.
+However, these degradations are small, and become smaller as the base model gets larger: corrected OPUS declines COMET, versus only for TI.
+These results also show the surprising difficulty of drop errors: although these are not more problematic for baseline models than other error types (see Tables 2 and 3), neither correction nor MT finetuning handles them as effectively as the other error types.
+This suggests that the missing information from a single deleted character often cannot be recovered from surrounding context.
+Effect of training size.
+Although we used 5 million training examples in our main experiments, Table 10 shows the effect that using fewer has on OPUS’s robustness.
+The correction approach performs much better when data is restricted to 100k examples, but this advantage shrinks as the training size is increased and eventually MT finetuning outperforms it.
+This supports the intuition that noisy translation is a harder task than correction, so the OPUS model requires more examples to learn it.
+How often does mitigation work?
+In the main results, we showed that MT finetuning and SC both greatly improve robustness.
+However, as the scores are presented at corpus level, they do not give insight into what percentage of examples are actually helped, versus how many are harmed.
+To fill in this missing piece, we compared sentence-level COMET between our baseline models and the two mitigation approaches.
+Results are shown in Figure 4.
+At high noise levels, both techniques nearly always help OPUS.
+However, they behave differently on clean sources.
+While SC almost never improves clean scores and harms less than 10% of examples, finetuning is more of a high-risk strategy, as it makes scores worse for about half of examples while also improving COMET for 38% of them.
+The trend is subtly different for SC with TI.
+While correction helps more examples than it hurts whenever the noise level is at least 10%, the percentage of examples that are harmed by correction actually increases up to 30% noise.
+This suggests that correction may introduce errors into some noisy sequences that TI could have handled itself.
+5.2 Mitigating Errors in MTNT
+Having shown that MT finetuning and SC are both effective techniques for improving robustness to synthetic errors, we return to MTNT.
+Our goal is to determine whether either MT finetuning or SC can provide some benefit to performance on enfr.
+| Finetuning | Uncorrected | Corrected |
+| None | 77.21 | 76.31 |
+| MTNT | 79.22 | 77.97 |
+| Synthetic | 75.95 | 75.31 |
+Finetuning vs. correction for OPUS.
+We evaluate MT finetuning and SC on MTNT.
+For MT finetuning, we compare two techniques: finetuning only on MTNT, as in Section 4.1; and finetuning on 5 million synthetically noised Tatoeba Challenge enfr examples, matching the procedure used for enpt in Section 5.1.
+For SC, we use the same English model as in the synthetic experiments.
+Results are shown in Table 11.
+Finetuning on MTNT outperforms the baseline, but both SC and MT finetuning are harmful, confirming that MTNT is not very noisy in terms of spelling errors (Karpukhin et al., 2019; Michel and Neubig, 2018).
+Effect of source correction.
+In Table 12, we show results with and without SC for various models.
+Although correction does not improve results on the whole, the oracle consistently outperforms the baseline by about COMET, illustrating that many sequences do benefit from correction.
+Indeed, 21.8% of MTNT examples improve with correction and another 46.1% are not harmed.
+Future work could investigate the use of a routing mechanism that decides for each example whether to apply SC.
+| Method | r/OPUS | NLLB | TI |
+| Base | 79.22 | 79.33 | 81.91 |
+| Correction | 77.97 | 78.32 | 80.94 |
+| Oracle | 79.71 | 79.82 | 82.43 |
+6 Conclusion
+We presented several experiments testing the robustness of MT systems to synthetic and natural noise.
+On synthetic noise, we showed that large multilingual MT models and LLMs are far more robust than older models.
+The experiments on social media translation showed that larger models also worked better on natural noise.
+We further supported this conclusion through reference-free translation experiments with a novel evaluation metric based on quality estimation.
+Finally, we showed that noisy finetuning and source correction allow smaller models to exceed GPT-3.5’s robustness synthetic noise, while also being useful in many cases for handling natural errors.
+Limitations
+We acknowledge the limitations of our work.
+All of the languages included in this study have large speaker populations and many resources available, and are the official languages of countries.
+Some conclusions may not generalize to low resource languages.
+Additionally, this paper studies only one source of natural noise, namely social media text.
+Other varieties of text perceived as noisy, such as transcribed speech or text written by nonfluent language users, may have different properties.
+Finally, the evaluation techniques used in this work are all automatic or neural, and may differ from gold-standard human evaluation.
+Acknowledgments
+This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
+We thank Sweta Agrawal and the anonymous reviewers for their helpful comments.
+Appendix A Key Noise
+As different languages customarily use different keyboard layouts, we made slight alterations to our key noising procedure for each source language.
+We use the QWERTZ layout for German, AZERTY for French, QWERTY for English and Portuguese, and South Korean Dubeolsik for Korean.
+For Korean, we used hangul-jamo999https://github.com/jonghwanhyeon/hangul-jamo to decompose hangul characters into jamo, which represent individual keystrokes, before applying perturbations.
+Appendix B OPUS Models
+| LP | Vocab | Params |
+| ende | 58.1k | 74.4M |
+| deen | 58.1k | 74.4M |
+| enes | 55.0k | 234.8M |
+| esen | 65.0k | 77.9M |
+| enfr | 53.0k | 232.7M |
+| fren | 53.0k | 232.7M |
+| enko | 32.0k* | 209.2M |
+| koen | 32.0k* | 209.2M |
+| enpt | 54.7k | 234.5M |
+| pten | 60.0k | 75.4M |
+The parameters and vocabulary sizes of the OPUS models are shown in Table 13.
+All checkpoints come from the Tatoeba Challenge (Tiedemann, 2020).101010https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/models
+Appendix C Inference
+Owing to the differing frameworks between OPUS, OPUSLLM, NLLB, TI, and ByT5 models, we use different beam search implementations depending on the model.
+For OPUS models, we decode with Marian (Junczys-Dowmunt et al., 2018), while NLLB and ByT5 use Hugging Face transformers.111111https://github.com/huggingface/transformers TI models and OPUSLLM are decoded with vllm (Kwon et al., 2023).121212https://github.com/vllm-project/vllm Regardless of framework, we used a beam size of across all experiments.
+C.1 Prompt for TowerInstruct
+‘‘Translate the following text from [source language] to [target language].nSource:[source text]n[target language]:’’
+C.2 Prompt for GPT-3.5
+‘‘Translate this sentence from [source language] to [target language].nSource:[source text]nTarget:’’
+Appendix D OPUSLLM Training Details
+The OPUSLLM model was trained with used Megatron-DeepSpeed131313https://github.com/microsoft/Megatron-DeepSpeed for a maximum of 300k steps with an effective batch size of 65k tokens and a base learning rate of , with a constant learning rate schedule and 5000 warmup steps.
+The model used the same 32k tokenizer as TI.
+To account for the model’s decoder-only structure, examples were formatted with two special tokens to indicate the beginning of the source and target sequences.
+At inference time, we selected the checkpoint with the best COMET on the FLORES enpt dev set.
+Appendix E Training Hyperparameters
+In this work, we finetune two types of base models: OPUS models and ByT5-Small.
+Despite the various technical differences between these two models, we used almost identical training procedures for them, with differences noted below.
+Basic training procedure.
+We finetuned using early stopping with patience 3. We validated every 500 steps. We used a grid over the learning rates and selected the best checkpoint by validation loss.
+Training library.
+For OPUS models, we finetuned with the marian command line tool from Marian (Junczys-Dowmunt et al., 2018).
+For ByT5-Small, we used a script that leverages the Trainer class from Hugging Face transformers.
+Appendix F MultiLexNorm Statistics
+Appendix G Intrinsic Correction Performance
+| swap | drop | dupe | key | |
 | None | 40.0 | 47.0 | 68.2 | 45.7 |
 | JamSpell | 75.5 | 63.5 | 91.7 | 78.0 |
-mT5-Large | 85.8 | 83.2 | 95.0 | 87.9 |
-We use two architectures for source correction:
-JamSpell (Ozinov, 2019) is a non-neural spell-checker with trigram context. We use the default English model from the package.171717https://github.com/bakwc/JamSpell
-We finetune mT5-based (Xue et al., 2021) correctors on the Prob+Word dataset from NeuSpell181818https://github.com/neuspell/neuspell (Jayanthi et al., 2020).
-We use the mT5-Large (1.2B parameter) version of the architecture with the hyperparameter ranges in Table 10.
-We select the checkpoint with the best chrF (Popović, 2015) when decoding a noised version of the FLORES English dev set.
-We decode with a beam size of 5.
-These two models are minimalistic and maximalistic approaches to correction.
-JamSpell is fast (more than 10k tokens per second on a laptop CPU) but uses only trigram context.
-On the other hand, mT5 is a large encoder-decoder that leverages a long context.
-Its subword vocabulary allows it to make open-vocabulary corrections, but it is much slower than JamSpell and is prone to hallucination.
-5.2 Correcting Synthetic Errors
-Model | ende | enfr | enko | enpt |
-OPUS | -0.44 | -0.45 | -0.30 | -0.38 |
-NLLB | -0.33 | -0.39 | -0.16 | -0.37 |
-Tower | -0.28 | -0.30 | 0.00 | -0.26 |
-GPT-3.5 | -0.30 | 0.32 | -0.10 | -0.32 |
-Model | ende | enfr | enko | enpt |
-OPUS | 0.10 | 0.00 | 0.01 | -0.01 |
-NLLB | -0.04 | -0.03 | -0.03 | 0.03 |
-Tower | -0.04 | -0.08 | -0.03 | -0.07 |
-GPT-3.5 | -0.11 | -0.08 | 0.07 | -0.04 |
-MT results.
-The effect of correcting synthetic enfr errors is shown in Figure 4.
-For OPUS models, source correction is often beneficial, with gains of more than 40 COMET points for mT5 over the baseline when at least 50% of tokens are noised.
-Source correction brings smaller performance gains for NLLB and Tower.
-Source correction has a slight negative effect on GPT-3.5, indicating that extremely robust models have little to gain from source correction.
-However, Table 12 shows that correcting clean inputs is nearly harmless.
-mT5 never loses more than COMET points.
-Even the weaker JamSpell never causes a degradation of more than half a COMET point, suggesting that it could be useful because of its fast runtime.
-5.3 Correcting MTNT
-Method | OPUS | NLLB | Tower | GPT-3.5 |
-Base | 77.51 | 79.20 | 80.65 | 81.33 |
-Finetuned | 79.37 | 80.85 | - | - |
-JamSpell | 76.42 | 78.09 | 79.17 | 79.68 |
-mT5 | 76.41 | 77.88 | 79.34 | 80.16 |
-Oracle | 78.65 | 80.11 | 81.59 | 82.16 |
-Results.
-Social media data is a challenge for both of our source correctors because they were trained on other domains.
-Using either corrected source leads to a decline of at least one COMET point compared to the raw source.
-Although an oracle delivers some gains, it still cannot equal the performance of finetuning when it is available.
-These results seem to confirm that MTNT is not particularly noisy in terms of spelling errors (Karpukhin et al., 2019; Michel and Neubig, 2018).
-Per-sentence analysis.
-Based on the results in Table 13, it would be easy to conclude that source correction does not work for MTNT.
-However, Figure 5 shows that some sentences do benefit from correction, and others are not harmed.
-In these terms, mT5 appears to be a higher-risk corrector than JamSpell, improving translation quality for of examples (versus for JamSpell) while making quality worse for of examples (versus for JamSpell).
-This, in addition to the oracle performance, suggests that further research could make source correction a useful tool in this domain.
-6 Conclusion
-We presented several experiments testing the robustness of MT systems to synthetic and natural noise.
-On synthetic noise, we showed that large multilingual MT models and LLMs are far more robust than older techniques.
-The experiments on social media translation showed that the same larger models also worked better on natural noise.
-We added further support for this conclusion through reference-free translation experiments with a novel evaluation metric based on quality estimation.
-Finally, we exhibited circumstances in which pipeline-based source correction techniques can improve performance on noisy text, both synthetic and natural.
-Acknowledgments
-This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882- 00000055 (Center for Responsible AI), and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020.
+| ByT5-Small | 97.5 | 89.6 | 99.6 | 94.4 |
+Intrinsic correction results in terms of chrF (Popović, 2015) are shown in Table 15.
+In addition to our ByT5-Small corrector, we include the results of the pretrained English model from JamSpell141414https://github.com/bakwc/JamSpell (Ozinov, 2019), a trigram-based spell-checker.
+Appendix H Synthetic Results with Other Metrics
+The models’ synthetic performance on clean data is shown in terms of BLEU151515nrefs:1|case:mixed|eff:no|tok:flores200|
+smooth:exp|version:2.5.1 in Table 16 and in chrF161616nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|
+version:2.5.1 in Table 17.
+Synthetic results showing BLEU-slope are shown in Tables 18 and 19.
+Synthetic results showing chrF-slope are shown in Tables 20 and 21.
+BLEU-slope and chrF-slope are computed analogously to COMET-slope.
+xxen
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+44.7
+49.7
+30.7
+49.3
+43.6
+NLLB
+49.3
+50.3
+34.8
+55.1
+47.4
+TI
+50.6
+52.1
+37.0
+56.7
+49.1
+GPT-3.5
+48.8
+49.1
+32.5
+53.6
+46.0
+enxx
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+42.0
+55.6
+23.8
+54.6
+44.0
+NLLB
+46.8
+56.2
+24.8
+54.5
+45.6
+TI
+47.0
+57.1
+29.5
+52.8
+46.6
+GPT-3.5
+47.8
+56.6
+25.6
+56.6
+46.7
+xxen
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+66.4
+69.1
+55.8
+69.4
+65.2
+NLLB
+69.2
+69.9
+58.6
+72.9
+67.7
+TI
+69.9
+70.6
+60.3
+73.9
+68.7
+GPT-3.5
+69.8
+69.7
+58.2
+72.7
+67.6
+enxx
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+62.9
+71.7
+36.2
+71.3
+60.5
+NLLB
+65.7
+71.8
+36.7
+71.0
+61.3
+TI
+66.2
+72.3
+38.7
+70.3
+61.9
+GPT-3.5
+67.1
+72.6
+35.1
+72.8
+61.9
+swap
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-51.4
+-58.2
+-28.2
+-56.2
+-48.5
+NLLB
+-33.4
+-35.8
+-30.6
+-37.5
+-34.3
+TI
+-37.9
+-38.8
+-30.4
+-39.6
+-36.7
+GPT-3.5
+-10.9
+-14.4
+-27.4
+-17.2
+-17.5
+drop
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-45.6
+-49.6
+-26.4
+-49.1
+-42.7
+NLLB
+-29.8
+-29.9
+-26.1
+-32.7
+-29.6
+TI
+-29.5
+-29.7
+-25.6
+-30.0
+-28.7
+GPT-3.5
+-14.4
+-13.8
+-24.9
+-16.0
+-17.3
+dupe
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-34.3
+-38.0
+-8.8
+-37.6
+-29.7
+NLLB
+-13.9
+-15.7
+-11.5
+-14.8
+-14.0
+TI
+-9.9
+-13.0
+-5.5
+-10.5
+-9.7
+GPT-3.5
+-5.5
+-5.3
+-12.0
+-5.8
+-7.2
+key
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-50.3
+-57.3
+-27.8
+-57.2
+-48.2
+NLLB
+-35.2
+-36.4
+-33.3
+-40.3
+-36.3
+TI
+-38.5
+-42.7
+-27.2
+-46.3
+-38.7
+GPT-3.5
+-18.3
+-18.2
+-22.5
+-21.9
+-20.2
+swap
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-49.1
+-62.6
+-28.9
+-64.5
+-51.3
+NLLB
+-31.2
+-33.8
+-18.1
+-37.1
+-30.0
+TI
+-19.8
+-20.7
+-12.6
+-20.0
+-18.3
+GPT-3.5
+-7.8
+-9.3
+-7.1
+-5.1
+-7.3
+drop
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-47.8
+-59.4
+-27.9
+-61.4
+-49.1
+NLLB
+-30.7
+-33.7
+-16.6
+-36.2
+-29.3
+TI
+-21.9
+-23.5
+-13.7
+-22.6
+-20.4
+GPT-3.5
+-12.0
+-13.4
+-9.6
+-12.8
+-12.0
+dupe
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-39.5
+-45.9
+-26.5
+-54.5
+-41.6
+NLLB
+-9.3
+-8.9
+-3.6
+-14.7
+-9.1
+TI
+-4.6
+-5.0
+-3.5
+-5.2
+-4.6
+GPT-3.5
+-3.7
+-4.0
+-3.2
+-3.1
+-3.5
+key
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-48.8
+-62.9
+-29.0
+-64.6
+-51.3
+NLLB
+-34.1
+-37.4
+-20.1
+-40.3
+-33.0
+TI
+-25.0
+-26.9
+-15.3
+-26.8
+-23.5
+GPT-3.5
+-11.8
+-13.1
+-10.5
+-12.5
+-12.0
+swap
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-45.2
+-47.4
+-32.9
+-46.0
+-42.9
+NLLB
+-25.3
+-26.6
+-27.7
+-26.4
+-26.5
+TI
+-28.4
+-27.3
+-29.0
+-28.1
+-28.2
+GPT-3.5
+-7.1
+-9.6
+-29.2
+-10.9
+-14.2
+drop
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-40.1
+-39.7
+-29.0
+-38.7
+-36.9
+NLLB
+-23.8
+-23.6
+-26.6
+-24.1
+-24.5
+TI
+-22.3
+-20.9
+-25.5
+-21.1
+-22.5
+GPT-3.5
+-10.0
+-9.3
+-26.0
+-10.4
+-13.9
+dupe
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-22.9
+-24.5
+-6.8
+-23.1
+-19.3
+NLLB
+-7.6
+-8.6
+-6.1
+-7.5
+-7.4
+TI
+-5.1
+-6.4
+-3.9
+-5.3
+-5.2
+GPT-3.5
+-2.7
+-2.7
+-10.2
+-2.8
+-4.6
+key
+Model
+deen
+fren
+koen
+pten
+avg.
+OPUS
+-43.6
+-45.5
+-34.5
+-46.0
+-42.4
+NLLB
+-26.6
+-27.7
+-31.0
+-29.6
+-28.7
+TI
+-28.6
+-30.6
+-26.3
+-34.6
+-30.0
+GPT-3.5
+-12.1
+-11.9
+-22.6
+-14.0
+-15.2
+swap
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-50.0
+-56.8
+-38.0
+-55.7
+-50.1
+NLLB
+-24.7
+-26.4
+-21.0
+-27.4
+-24.9
+TI
+-15.9
+-15.0
+-11.8
+-14.8
+-14.4
+GPT-3.5
+-5.1
+-6.0
+-6.4
+-5.1
+-5.7
+drop
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-48.3
+-53.3
+-36.6
+-53.5
+-47.9
+NLLB
+-24.7
+-26.6
+-20.3
+-27.2
+-24.7
+TI
+-17.4
+-17.0
+-12.7
+-16.6
+-15.9
+GPT-3.5
+-8.3
+-9.0
+-8.4
+-8.2
+-8.5
+dupe
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-34.7
+-36.8
+-33.6
+-41.8
+-36.7
+NLLB
+-5.7
+-5.8
+-5.7
+-8.6
+-6.4
+TI
+-2.8
+-2.6
+-2.8
+-2.7
+-2.7
+GPT-3.5
+-1.9
+-2.1
+-2.5
+-1.3
+-2.0
+key
+Model
+ende
+enfr
+enko
+enpt
+avg.
+OPUS
+-50.0
+-56.6
+-37.9
+-55.9
+-50.1
+NLLB
+-27.9
+-29.8
+-22.7
+-30.4
+-27.7
+TI
+-21.0
+-20.6
+-15.3
+-20.5
+-19.4
+GPT-3.5
+-8.4
+-9.0
+-8.8
+-8.1
+-8.6